OpenStudy (anonymous):

Find the exact value by using a half-angle identity. tan 7pi/8

zepdrix (zepdrix):

So we'll start from here. \[\huge \tan \left(\frac{7\pi}{8}\right)=\tan \left(\frac{\left(\frac{7\pi}{4}\right)}{2}\right)\]

zepdrix (zepdrix):

\[\huge \tan \left(\frac{a}{2}\right)=\frac{1-\cos a}{\sin a}\]Using this identity gives us,

zepdrix (zepdrix):

\[\huge \frac{1-\cos \left(\frac{7\pi}{4}\right)}{\sin \left(\frac{7\pi}{4}\right)}\]And from here, you just need to remember your special angles on the unit circle.

OpenStudy (anonymous):

and that is?

zepdrix (zepdrix):

cos(7pi/4) and the sine also, you need to remember those. If you don't know, go back to your chart or whatever you use to figure out the special angles :D It's one of the pi/4 angles, so it'll be something involving sqrt(2)/2's

OpenStudy (anonymous):

https://www.youtube.com/watch?v=Nh4Ro86dnHk